OpenStudy (anonymous):

Partial Fractions question:

OpenStudy (anonymous):

I realise the below equation is true but how would i have known this was true if normally with partial fractions you add the two terms on the right hand side of the equals. \[\large \frac{1}{i^{2}+3i+2} = \frac{1}{(i+2)(i+1)} = \frac{1}{i+1} - \frac{1}{i+2}\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

...but this time you minus them

OpenStudy (unklerhaukus):

\[ \frac{1}{i^{2}+3i+2} = \frac{1}{(i+2)(i+1)} = \frac{A}{i+1} + \frac{B}{i+2}\]

OpenStudy (unklerhaukus):

\[(i+2)(i+1)\times \frac{1}{(i+2)(i+1)} = (i+2)(i+1)\times\left(\frac{A}{i+1} + \frac{B}{i+2}\right)\] \[\qquad\qquad\qquad\qquad1= (i+2)A+(i+1)B\]

OpenStudy (unklerhaukus):

if \( i =-1\) \[1=A\] if \(i=-2\) \[1=-B\]

OpenStudy (unklerhaukus):

can you follow that?

OpenStudy (anonymous):

yeah....wasn't thinking it through properly. thanks :)

OpenStudy (amistre64):

yeah, Unkle did the proper steps which most teachers ive seen omit. is "i" used here as a normal variable, or is it the complex i ?

OpenStudy (amistre64):

not that it matter much, the complex i is treated as a variable until the end when it is simplified